OpenStudy (anonymous):

In acceleration vs time grapg of a Particle covering a distance of 40m from t=2 till t=4s The magnitude of Velocity at t= 10s

OpenStudy (anonymous):



OpenStudy (ajprincess):

I am nt sure if this is right. Using s=ut+1/2at^2 40=2u+1/2*5*4 40=2u+10 2u=40-10 2u=30 u=30/2 u=15 using v=u+at v=15+5*2 v=15+10 v=25 The area

OpenStudy (anonymous):

25 + 15 = 40

OpenStudy (anonymous):

right ?

OpenStudy (anonymous):

1/2 * 5 * 6 = 15

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

but @ajprincess i want to know why Area under this Graph results in 30 m/s

OpenStudy (ajprincess):

It is not given that the particle starts from rest. so it might have had an initial velocity. From the graph acc is only zero at first nt the velocity.

OpenStudy (ajprincess):

does that make sense @yahoo?

OpenStudy (anonymous):

What exactly are you supposed to find ?

OpenStudy (ajprincess):

The magnitude of Velocity at t= 10s

OpenStudy (anonymous):

thxxx

OpenStudy (ajprincess):

welcome:)